OpenStudy (brooke..help00):

PLEASE HELP :) Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. (5 points) Part B: How many liters of orange juice are present in the can? Show your work. __I know that .2(water in the can= 1liters) (so orange juice in the can .8=4 liters.) I am confused about the equation to write though. + @TKHUNNY @MATHMALE @WILL.H @HERO @RHONDASOMMER @KITTIWITTI1 @JIGGLYPUFF314 +

RhondaSommer (rhondasommer):

hm. 80% OJ 20% water \[\frac{20}{100} = \frac{1}{x}\] x=5 liters total 5-1(the liter of water) then = 4 (the amount of OJ. BTW WATERING DOWN OJ IS DISGUSTING

RhondaSommer (rhondasommer):

Also, I dont know if that is exactly the way your teacher wants it done.

OpenStudy (brooke..help00):

I can format it to the equation. Thanks so much I was getting confused. and LOL.